# Self-Evolving Multi-Agent Simulations for Realistic Clinical Interactions

### v1

@@ Abstract
In this work, we introduce MedAgentSim, an open-source simulated clinical environment with doctor, patient, and measurement agents designed to evaluate and enhance LLM performance in dynamic diagnostic settings. Unlike prior approaches, our framework requires doctor agents to actively engage with patients through multi-turn conversations, requesting relevant medical examinations (e.g., temperature, blood pressure, ECG) and imaging results (e.g., MRI, X-ray) from a measurement agent to mimic the real-world diagnostic process. Additionally, we incorporate self improvement mechanisms that allow models to iteratively refine their diagnostic strategies. We enhance LLM performance in our simulated setting by integrating multi-agent discussions, chain-of-thought reasoning, and experience-based knowledge retrieval, facilitating progressive learning as doctor agents interact with more patients. We also introduce an evaluation benchmark for assessing the LLM’s ability to engage in dynamic, context-aware diagnostic interactions. While MedAgentSim is fully automated, it also supports a user-controlled mode, enabling human interaction with either the doctor or patient agent. Comprehensive evaluations in various simulated diagnostic scenarios demonstrate the effectiveness of our approach. Our code, simulation tool, and benchmark are available on the project page.
Keywords:
Multi Agents Visual Agents Self Improving Agents.
**footnotetext: These authors contributed equally to this work.
@@ 1 Introduction
Advancements in Large Language Models (LLMs) and Vision-Language Models (VLMs) have shown promising capabilities across various medical tasks, achieving human-level performance on several medical benchmarks [32]. These models have demonstrated the ability to encode clinical knowledge [41, 47], retrieve relevant medical literature [53], and achieve high accuracy in single-turn medical question-answering tasks [7, 22, 32, 51]. However, current medical LLM assessments often rely on static evaluation benchmarks, where models are provided with complete patient information and tasked with answering predefined questions, sometimes with multiple-choice options [17]. These assessments often fail to capture the complexity of real-world doctor-patient interactions, where diagnosis is not a single-step process but a dynamic, multi-turn dialogue [19]. Such multi-turn doctor-patient interactions are important in clinical scenarios, as patients often struggle to describe their symptoms accurately due to limited medical knowledge, ambiguous perceptions, or communication barriers [29]. Consequently, physicians play an active role in structuring these interactions, posing clarifying questions, and refining their assessments as new information emerges [60].
@@
    A multi-agent diagnostic pipeline that improves baseline LLM performance by incorporating self-improvement mechanisms, including multi-agent discussion, chain-of-thought reasoning, and experience-based knowledge retrieval. The system enables progressive learning, where doctor agents refine their diagnostic capabilities as they interact with more patients.
2 Related Work
LLMs in the Medical Field The application of Large Language Models (LLMs) in medicine has been extensively explored, with early efforts focusing on domain-specific pretraining to enhance performance on biomedical tasks. This approach has proven effective, as demonstrated by models such as PubMedBERT [14] and BioGPT [26], which leverage self-supervised objectives trained on specialized corpora. Several other models, including BioLinkBERT [58], BioMedX [31], DRAGON [57], BioMedLM [5, 9], and MedPaLM [40], have followed similar strategies, refining their capabilities for medical question answering, health inquiries, and doctor-patient dialogues. Additionally, fine-tuned models such as DoctorGLM [54], Bianque2 [6], ChatMed-Consult [61], MedicalGPT [55], and DISCMedLLM [3] have been developed using diverse datasets and optimization frameworks to improve model adaptability to medical contexts.
More recently, prompt-based approaches have emerged as a competitive alternative to domain-specific pretraining and fine-tuning. Methods such as MedPrompt [32], OpenMedLM [27], Prompt-Eng [1], and related works [4] leverage foundation models without requiring additional pretraining, instead relying on carefully designed prompt engineering techniques to achieve state-of-the-art results in medical question answering. These works demonstrate that prompt engineering alone can outperform fine-tuning strategies in certain medical tasks, highlighting the efficiency and adaptability of foundation models in the healthcare domain.
Multi-Agent LLMs in the Medical Field The multi-agent paradigm in LLM research has gained significant traction, leveraging the planning and reasoning capabilities of LLMs as autonomous agents for complex problem-solving [10, 15, 42, 59]. While single-agent LLMs have shown proficiency in tasks such as decision-making and diagnostic assistance [24, 32], recent efforts have explored multi-agent frameworks where multiple LLMs collaborate, mirroring the division of labor seen in real-world medical practice [13, 39, 49].
These multi-agent setups aim to address the limitations of single-agent decision-making by introducing specialized agents, each trained or prompted to handle specific aspects of a medical task [43, 50]. By incorporating group collaboration structures, these systems optimize complex workflows, ensuring more accurate and contextually aware medical assessments. Notably, MedAgents [43] demonstrated that multi-agent architectures outperform single-agent LLMs in medical reasoning tasks, particularly in handling specialized diagnoses and treatment planning. Such findings reinforce the importance of agent collaboration in medical AI applications, offering scalable and modular approaches for improving LLM-based healthcare solutions.
Simulated Agents in the Medical Field Most LLM-based medical AI systems operate in a static question-answering format, where models are provided full patient details upfront and expected to generate responses from a predefined set of multiple-choice answers [1, 4, 27, 32]. While this approach is useful for benchmarking, it fails to capture the real-world complexities of patient-doctor interactions, where information is often incomplete, and medical practitioners must dynamically query patients to obtain necessary data [13].
To address this limitation, recent research has focused on simulating realistic patient-doctor interactions using LLM agents [21]. Early work in LLM-powered simulations demonstrated that when placed in interactive environments with structured memory and reasoning capabilities, LLMs could exhibit human-like behavior in decision-making processes [30, 36, 52]. Building upon this, efforts have been made to simulate AI-driven hospitals, where LLM-powered doctors and patients engage in evolving interactions to improve diagnostic accuracy [11, 21]. These simulations enable agents to learn from their mistakes mid-simulation, refining their diagnostic reasoning over multiple patient interactions, similar to how human doctors develop expertise through experience [11, 21].
By embedding medical LLM agents within a game-based simulation, researchers have introduced a new paradigm for evaluating and training AI-driven diagnostic models. Unlike traditional static evaluations, these dynamic simulations allow LLM agents to iteratively improve, adapt to evolving medical scenarios, and enhance their performance through experience-driven learning. This simulation-based framework presents a promising avenue for developing autonomous AI medical assistants capable of operating in realistic, interactive healthcare environments.
@@ 3 Methodology: MedAgentSim
(a) In Conversation Phase, the doctor and patient agents engage in an interactive dialogue, allowing the doctor to gather vital information and request necessary diagnostic tests, such as blood tests and X-Rays, from the measurement agent. As results are provided, the conversation continues until the doctor has sufficient information. Once ready to diagnose, the process transitions to Experience Replay Phase. Here, past doctor-patient interactions are analyzed through memory buffers, retrieving relevant cases as few-shot examples to enrich the current dialogue. A team of doctor agents then evaluates this enhanced conversation using chain-of-thought reasoning and majority-vote ensembling to reach a consensus, producing a well-informed diagnosis.
@@
(a) Agent Roles. The simulation consists of three core agent types: the patient agent, the doctor agent, and the measurement agent. The patient experiences symptoms and seeks medical attention from the doctor, who is responsible for diagnosing and treating conditions. The measurement agent provides diagnostic test results but only when explicitly requested, requiring the doctor to actively gather information rather than receiving all patient data upfront. Figure 1(b) showcases a sample scenario, demonstrating how agents navigate the environment and engage in clinical workflows. This baseline framework is referred to as Multi-Agent Clinic.This framework is referred to as our baseline.
@@
(a) Records Buffer. To enable progressive learning, the system maintains a record storage and retrieval mechanism that captures both successful and corrected diagnoses. It consists of two dynamically expanding libraries: the Medical Records Buffer, which stores correctly diagnosed cases, and the Experience Records Buffer, which retains misdiagnosed cases that were later corrected through reflection. During a new consultation, the system uses k-nearest neighbors (KNN) to retrieve relevant past cases. The Medical Records Buffer provides full conversations and diagnoses, while the Experience Records Buffer extracts key insights from the reflection process. This approach leverages prior experiences, as studies show that LLMs improve accuracy when learning from failures [56].
(b) COT and Ensembling. The retrieved information is then incorporated into the consultation, enriching the doctor’s contextual understanding. A multi-agent system processes the updated input, where multiple doctor agents independently assess the case and propose diagnoses. These assessments are aggregated and refined using chain-of-thought reasoning and majority-vote ensembling [32], producing a final diagnosis.
(c) Records Storage. Once finalized, the system converts each component of the case, such as conversation history, diagnosis, medical images, and lab results, into CLIP [38] embeddings. For lengthy inputs like conversation history, an LLM (the same one powering the agents) first summarizes the content into 2–3 sentences. When a diagnosis embedding is matched, associated embeddings from the same case (e.g., conversation history or lab results) are retrieved alongside it, enabling contextual grounding. Correct diagnosis embeddings are added to the Medical Records Buffer, while incorrect cases trigger a reflection phase in which the doctor analyzes the mistake before making a second attempt. If the revised diagnosis is correct, only the CLIP-embedded reflection insights are stored in the Experience Records Buffer; otherwise, the case is discarded to ensure learning is based on meaningful examples. Figure 2(b) illustrates the full reflection and storage process.
@@ 4 Experiments
Experimental Details. We conducted extensive experiments to evaluate the effectiveness of MedAgentSim in a real-world doctor-agent setting. Our study leveraged a diverse set of both open-source models available on Hugging Face [8] and proprietary models, tested across three primary benchmarks: NEJM [39], MedQA [17], and MIMIC-IV [18].
For VLM tasks, we utilized the NEJM dataset, which includes 15 complex real-world cases along with an extended set, NEJM Extended, of 120 additional cases. MedQA comprises 106 simulated diagnostic scenarios, while its extended variant, MedQA Extended, contains 214 distinct cases. Additionally, MIMIC-IV features 288 clinical cases, providing a diverse set of real-world medical interactions. As these datasets are primarily formatted for QA tasks, they are not directly compatible with our simulation pipeline. To address this, we preprocess the data using GPT-4o, converting it into a structured JSON format, where the doctor, patient information, and test results are assigned to the doctor agent, patient agent, and measurement agent, respectively. Model accuracy is evaluated using a binary true/false metric for the final diagnosis, with an LLM serving as the evaluator to account for variability in generated responses. Both the dataset conversion process and accuracy logs were manually reviewed to ensure reliability.
@@
Results and Analysis. Table 1 compares the performance of the baseline Multi-Agent Clinic and our proposed MedAgentSim across key medical benchmarks, covering both language-based and vision-based tasks. We follow a similar evaluation strategy to samuel et al. [39] across all the benchmarks. MedAgentSim integrates LLaVA 1.5-Mistral, a multimodal model combining visual encoding with large language models.
@@
Table 1: Performance of Multi-Agent Clinic and MedAgentSim (Our) models across medical benchmarks. We used diverse LLMs including closed source. For visual language tasks, we use LLava 1.5 [25] for visual encoding.
@@ 4.1 Ablation Study
Impact of MedAgentSim Strategies. Table 2 summarizes the impact of adding incremental reasoning strategies on model accuracy. The integration of measurement, memory augmentation, chain-of-thought (COT) reasoning, and ensembling progressively improves diagnostic performance. Notably, the LLaMa 3.3 70B model benefits significantly from memory and COT strategies, achieving a final accuracy boost of 16.1%.
Table 2: Incremental improvements in model accuracy as measurement, memory, COT, and ensembling techniques are added.
@@
5 Ethical Considerations
This work involves the simulation of clinical environments using large language models (LLMs) and synthetic patient data. As MedAgentSim simulates diagnostic interactions and decision-making processes, it is important to emphasize that the system is intended strictly for research purposes and is not designed or validated for real-world clinical use. Any deployment of similar AI systems in healthcare settings must undergo rigorous clinical validation, regulatory approval, and expert oversight.
We acknowledge the potential for bias in LLM outputs, particularly when models are trained on large-scale web data that may encode societal and medical biases. To address this, we conducted a bias analysis to evaluate disparities in model performance and highlight the importance of fairness-aware model development. However, further research is necessary to ensure equitable and safe AI behavior across diverse populations and medical contexts.
Finally, we advocate for transparency and reproducibility in AI research. To that end, MedAgentSim is released as an open-source framework, allowing the broader community to audit, extend, and build upon this work while promoting responsible AI development in healthcare.
@@ 6 Conclusion
We introduced MedAgentSim, a multi-agent framework for interactive doctor-patient simulations that enhances diagnostic accuracy through structured reasoning, measurement-based decision-making, and self-improvement mechanisms. Our results demonstrate that memory, chain-of-thought prompting, and ensembling significantly improve performance in realistic clinical scenarios. Additionally, our bias analysis highlights disparities in model robustness, emphasizing the need for fairness-aware AI in clinical applications. By bridging the gap between static benchmarks and real-world diagnostic reasoning, MedAgentSim provides a more adaptive approach to AI-driven healthcare.

### v2

# Equal contribution  Corresponding author

@@ Abstract
In this work, we introduce MedAgentSim, an open-source simulated clinical environment with doctor, patient, and measurement agents designed to evaluate and enhance LLM performance in dynamic diagnostic settings. Unlike prior approaches, our framework requires doctor agents to actively engage with patients through multi-turn conversations, requesting relevant medical examinations (e.g., temperature, blood pressure, ECG) and imaging results (e.g., MRI, X-ray) from a measurement agent to mimic the real-world diagnostic process. Additionally, we incorporate self improvement mechanisms that allow models to iteratively refine their diagnostic strategies. We enhance LLM performance in our simulated setting by integrating multi-agent discussions, chain-of-thought reasoning, and experience-based knowledge retrieval, facilitating progressive learning as doctor agents interact with more patients. We also introduce an evaluation benchmark for assessing the LLM’s ability to engage in dynamic, context-aware diagnostic interactions. While MedAgentSim is fully automated, it also supports a user-controlled mode, enabling human interaction with either the doctor or patient agent. Comprehensive evaluations in various simulated diagnostic scenarios demonstrate the effectiveness of our approach. Our codebase, simulation environment, and benchmark datasets are publicly available on the project page: https://medagentsim.netlify.app/.
@@ 1 Introduction
Advancements in Large Language Models (LLMs) and Vision-Language Models (VLMs) have shown promising capabilities across various medical tasks, achieving human-level performance on several medical benchmarks [17]. These models have demonstrated the ability to encode clinical knowledge [25, 29], retrieve relevant medical literature [33], and achieve high accuracy in single-turn medical question-answering tasks [2, 12, 17, 32]. However, current medical LLM assessments often rely on static evaluation benchmarks, where models are provided with complete patient information and tasked with answering predefined questions, sometimes with multiple-choice options [8]. These assessments often fail to capture the complexity of real-world doctor-patient interactions, where diagnosis is not a single-step process but a dynamic, multi-turn dialogue. Such multi-turn doctor-patient interactions are important in clinical scenarios, as patients often struggle to describe their symptoms accurately due to limited medical knowledge, ambiguous perceptions, or communication barriers [16]. Consequently, physicians play an active role in structuring these interactions, posing clarifying questions, and refining their assessments as new information emerges [35].
@@
    A multi-agent diagnostic pipeline that improves baseline LLM performance by incorporating self-improvement mechanisms, including multi-agent discussion, chain-of-thought (COT) [31] reasoning, and experience-based knowledge retrieval. The system enables progressive learning, where doctor agents refine their diagnostic capabilities as they interact with more patients.
@@ 2 Methodology: MedAgentSim
(a) In Conversation Phase, the doctor and patient agents engage in an interactive dialogue, allowing the doctor to gather vital information and request necessary diagnostic tests, such as blood tests and X-Rays, from the measurement agent. As results are provided, the conversation continues until the doctor has sufficient information. Once ready to diagnose, the process transitions to Experience Replay Phase. Here, past doctor-patient interactions are analyzed through memory buffers, retrieving relevant cases as few-shot examples to enrich the current dialogue. A team of doctor agents then evaluates this enhanced conversation using COT reasoning and majority-vote ensembling to reach a consensus, producing a well-informed diagnosis.
@@
(a) Agent Roles. The simulation consists of three core agent types: the patient agent, the doctor agent, and the measurement agent. The patient experiences symptoms and seeks medical attention from the doctor, who is responsible for diagnosing and treating conditions. The measurement agent provides diagnostic test results but only when explicitly requested, requiring the doctor to actively gather information rather than receiving all patient data upfront. Figure 1(b) showcases a sample scenario, demonstrating how agents navigate the environment and engage in clinical workflows. This baseline framework is referred to as Multi-Agent Clinic [24].
@@
(a) Records Buffer. To enable progressive learning, the system maintains a record storage and retrieval mechanism that captures both successful and corrected diagnoses. It consists of two dynamically expanding libraries: the Medical Records Buffer, which stores correctly diagnosed cases, and the Experience Records Buffer, which retains misdiagnosed cases that were later corrected through reflection. During a new consultation, the system uses k-nearest neighbors (KNN) to retrieve relevant past cases. The Medical Records Buffer provides full conversations and diagnoses, while the Experience Records Buffer extracts key insights from the reflection process. This approach builds on prior experiences, as studies show that LLMs benefit from failure-driven learning [34].
(b) COT and Ensembling. The retrieved information is then incorporated into the consultation, enriching the doctor’s contextual understanding. A multi-agent system processes the updated input, where multiple doctor agents independently assess the case and propose diagnoses. These assessments are aggregated and refined using COT [31] reasoning and majority-vote ensembling [17], producing a final diagnosis.
(c) Records Storage. Once finalized, the system converts all case data, including conversation history, diagnosis, medical images, and lab results, into CLIP [23] embeddings. Correct diagnosis embeddings are added to the Medical Records Buffer, while incorrect cases undergo a reflection phase where the doctor analyzes the mistake before making a second attempt. If the revised diagnosis is correct, only the CLIP-embedded reflection insights are stored in the Experience Records Buffer; otherwise, the case is discarded to ensure learning is based on meaningful examples. Figure 2(b) illustrates the full reflection and storage process.
@@ 3 Experiments
Experimental Details. We conducted extensive experiments to evaluate the effectiveness of MedAgentSim in a real-world doctor-agent setting. Our study leveraged a diverse set of both open-source models available on Hugging Face [3] and proprietary models, tested across three primary benchmarks: NEJM [24], MedQA [8], and MIMIC-IV [9]. For VLM tasks, we utilized the NEJM dataset, which includes 15 complex real-world cases along with an extended set, NEJM Extended, of 120 additional cases. MedQA comprises 106 simulated diagnostic scenarios, while its extended variant, MedQA Extended, contains 214 distinct cases. Additionally, MIMIC-IV features 288 clinical cases, providing a diverse set of real-world medical interactions. As these datasets are primarily formatted for QA tasks, they are not directly compatible with our simulation pipeline. To address this, we preprocess the data using GPT-4o, converting it into a structured JSON format, where the doctor, patient information, and test results are assigned to the doctor agent, patient agent, and measurement agent, respectively. Model accuracy is evaluated using a binary true/false metric for the final diagnosis, with an LLM serving as the evaluator to account for variability in generated responses. Both the dataset conversion process and accuracy logs were manually reviewed to ensure reliability.
@@
Results and Analysis. Table 1 compares the performance of the baseline Multi-Agent Clinic and our proposed MedAgentSim across key medical benchmarks, covering both language-based and vision-based tasks. MedAgentSim integrates LLaVA 1.5-Mistral, a multi-modal model combining visual encoding with large language models.
@@
In addition to automated benchmarks, a preliminary human study was performed in which one agent (doctor, patient, or measurement) was replaced with a real human. Doctors viewing the clips failed to identify the human in 62.5% of cases, indicating high behavioral realism. Full results are on the project page [^1] https://medagentsim.netlify.app.
Table 1: Performance of Multi-Agent Clinic (Basic) and MedAgentSim (Our) models across medical benchmarks. We used diverse LLMs including closed source. For visual language tasks, we use LLava 1.5 [14] for visual encoding.
@@ 3.1 Ablation Study
Impact of MedAgentSim Strategies. Table 2 summarizes the impact of adding incremental reasoning strategies on model accuracy. The integration of measurement, memory augmentation, COT [31] reasoning, and ensembling progressively improves diagnostic performance. Notably, the LLaMa 3.3 70B model benefits significantly from memory and COT strategies, achieving a final accuracy boost of 16.1%.
Table 2: Incremental improvements in model accuracy as measurement, memory, COT [31], and ensembling techniques are added.
@@
Table 3: Doctors’ ability to identify the real human in each video. ✓ = Correct guess, ✗ = Incorrect guess.
Video ID Doctor ID Guess Ground Truth Accuracy
VIDEO_01 Doctor 1 None None (All AI)
Doctor 2 Patient
VIDEO_02 Doctor 1 Patient Doctor
Doctor 2 None
VIDEO_03 Doctor 1 Measurement Measurement
Doctor 2 Patient
VIDEO_04 Doctor 1 Doctor Patient
Doctor 2 Patient
Total Correct (Doctor 1) 2 / 4 (50%)
Total Correct (Doctor 2) 1 / 4 (25%)
Combined Accuracy 3 / 8 (37.5%)
Combined Failure Rate 5 / 8 (62.5%)
@@
Table 4: Summary of expert perception accuracy when evaluating simulated clinical interactions.
Evaluation Metric Result
Real humans misidentified as AI 66.7%
AI agents misidentified as human 41.7%
Overall accuracy of identifying human agents 37.5%
Overall misidentification rate (human or AI) 62.5%
@@ 4 Conclusion
We introduced MedAgentSim, a multi-agent framework for interactive doctor-patient simulations that enhances diagnostic accuracy through structured reasoning, measurement-based decision-making, and self-improvement mechanisms. Our results demonstrate that memory, COT prompting, and ensembling significantly improve performance in realistic clinical scenarios. Additionally, our bias analysis highlights disparities in model robustness, emphasizing the need for fairness-aware AI in clinical applications. By bridging the gap between static benchmarks and real-world diagnostic reasoning, MedAgentSim provides a more adaptive approach to AI-driven healthcare.
Acknowledgments. We gratefully acknowledge support for this work from the Meta Llama Impact Innovation Award, the Meta Regional Research grant (Project OMER), the Google research award, the NVIDIA Academic grant, and the MBZUAI-WIS research grant (P008).
Disclosure of Interests. The authors declare no competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.